# Utilize the Flow before Stepping into the Same River Twice: Certainty Represented Knowledge Flow for Refusal-Aware Instruction Tuning

### v2

@@ 4.1 Overview
Figure 6: The Framework of CRaFT: Stage 1 queries the knowledge state and flow, while Stage 2 constructs and tunes refusal-aware instructions.
@@ 4.2 Query the Knowledge State and Flow of LLM
Correctness is obtained through exact match across the $N$ responses, calculating the proportion of accurate answers. Certainty is evaluated using a pretrained SentenceTransformer model  [^6] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to encode each response $\hat{a}_{j}$ into its vector representation $E(\hat{a}_{j})$, and the average similarity is computed between these vector embeddings (excluding diagonal elements). The correctness values range from $[0,1]$. In MCQA task, certainty ranges from [$-Entropy(N_{o})$, 0], and for OEQA, from $[0,1]$. $N_{o}$ stands for the number of options. More details about knowledge state query are in Appendix B.1.
@@
4.3 Refusal-Aware Instruct Data Constuction
In contrast to Cor-RAIT, which selects RAIT samples solely based on correctness, our approach leverages four parameters $\mu$, $\sigma$, $\Delta\mu$, and $\Delta\sigma$ to characterize both the knowledge state and flow of $\mathcal{M}_{0}$. The challenge lies in making informed sample selections across these four dimensions. We propose a two-step heuristic method outlined in Algorithm 1.
@@ 5.1 Dataset
In this paper, we evaluate two tasks: knowledge-oriented Multiple Choice Questions Answer (MCQA) and Open-ended Questions Answer (OEQA). For the MCQA task, the MMLU (Hendrycks et al. 2020) test split is used as the training set, the MMLU val split is used as the In-Domain (ID) test set, and the ARC-c (Clark et al. 2018) test split is the Out-Of-Domain (OOD) test set. For the OEQA task, the TriviaQA (Joshi et al. 2017) train split is used as the training set, the TriviaQA dev split is the ID test set, and the NQ (Kwiatkowski et al. 2019) dev split is the OOD test set. Refer to Appendix D.1 for details of the dataset.
@@ 5.2 Metric
In the post-RAIT evaluation of LLMs, each test sample is classified as correct, incorrect, or refused. We calculate the accuracy ($P_{c}$), error ($P_{w}$), and refusal rates ($P_{r}$) based on these categories. These metrics reveal trade-offs in LLM performance, prompting the essential question: How to determine the best-performing model?
@@
We constructed the initial model $\mathcal{M}_{0}$ along with four refined models $\mathcal{M}_{1}$ to $\mathcal{M}_{4}$, with exhibiting exemplary properties. Performance comparisons clearly demonstrate a progressive improvement: $\mathcal{M}_{1}<\mathcal{M}_{2}<\mathcal{M}_{3}<\mathcal{M}_{4}$. Detailed information on $\mathcal{M}_{0}$ through $\mathcal{M}_{4}$ can be found in Appendix C.2. However, existing metrics exhibit significant issues: $S_{\text{honesty}}$ (Yang et al. 2023) inaccurately rates $\mathcal{M}_{1}$ higher than $\mathcal{M}_{2}$ and equates $\mathcal{M}_{3}$ with $\mathcal{M}_{4}$. TRUTHFUL (Cheng et al. 2024) mistakenly favor $\mathcal{M}_{1}$ over $\mathcal{M}_{2}$. R-Acc, ER, A-Acc and A-F1 (Feng et al. 2024) also incorrectly ranks $\mathcal{M}_{1}$ higher than $\mathcal{M}_{2}$. AP (Zhang et al. 2024) fails to differentiate between $\mathcal{M}_{3}$ and $\mathcal{M}_{4}$.
@@ 5.3 Baselines
To verify the effectiveness of CRaFT, we compared it with the current mainstream methods: Init-Basic: We use the initial LLM, based on common question-answering prompts to instruct the LLM to answer questions. Init-Refuse: Compared to Init-Basic, we add instructions like “If you do not know the answer, please respond with ‘I don’t know.’” in the prompt. Van-Tuning: We randomly select $N_{\text{van}}+N_{\text{idk}}$ samples from $D_{\text{src}}$ for instruct-tuning directly, without sample modification. Cor-RAIT: We implemented this method following the settings in (Zhang et al. 2024), which is based on the correctness of the reply to filter and modify RAIT data. The prompts involved in each baseline are detailed in the Appendix D.2.
@@ 5.4 Implementation Details
In the experiments, we selected LLaMA2-7B-Chat (Touvron et al. 2023) and LLaMA3-8B-Instruct (Dubey et al. 2024) as the the initial LLM $\mathcal{M}_{0}$. In the MCQA task, we selected 5000 samples from MMLU as training data; in the OEQA tasks, we selected 10,000 samples from TriviaQA as training data. Except for Van-Tuning, where all samples are vanilla samples, in other RAIT settings, the ratio of vanilla and IdK samples is 1:4. In MCQA and OEQA tasks 5-shot and 3-shot were used in knowledge state query respectively. More details of the LLM’s knowledge state and flow query are in Appendix B. We used XTuner  [^7] https://github.com/InternLM/xtuner for all Instruct Tuning experiments, with 3 epochs and a maximum context length of 2048. In the MCQA tasks, we used LoRA, setting $r=64$, $\alpha=16$, $dropout=0.1$, and a learning rate of 2e-4; in the OEQA tasks, we used full parameter training. More details of training are in Appendix D.3. 0-shot and greedy decoding setting were used in evaluation. More details of evaluation are in Appendix D.4. OpenCompass  [^8] https://github.com/open-compass/opencompass was used for all knowledge state queries and evaluations. All experiments were conducted on NVIDIA A100-80GB GPUs.
@@ 6.1 Overall Performance
The experimental results on the OEQA and MCQA tasks are presented in Table 2 and 3. Under the ID setting for both types of tasks, our method outperformed four baseline models on THS, achieving the best results. Specifically, under the ID setting for OEQA, compared to the current best RAIT baseline, CRaFT improved the THS on LLaMA2-7B-Chat and LLaMA3-8B-Instruct by 3.56 and 2.72, respectively. Similarly, under the ID setting for MCQA, CRaFT improved the THS by 1.14 and 13.57, respectively. This indicates that CRaFT can significantly improve the model’s rejection capability under the ID setting. Under the OOD setting, CRaFT improved the THS on the MCQA task by 1.5 and 9.2, respectively, compared to Cor-RAIT. On the OEQA’s LLaMA2-7B-Chat, it improved by 1.76 compared to the most competitive method, Init-Refuse. Overall, CRaFT demonstrated excellent competitiveness in model generalization. Furthermore, we found that on the MCQA task, compared to other baselines, Cor-RAIT showed significant improvements under both ID and OOD settings. However, on the OEQA task, Cor-RAIT performed worse than Init-Refuse under the OOD setting. This reveals the limitations of the instruction fine-tuning method. It’s worth mentioning that Van-Tuning generally had a negative impact on the improvement of overall capability, implying that the instruction fine-tuning approach of forcing the model to answer can undermine the model’s inherent rejection capability. Therefore, although CRaFT surpassed Cor-RAIT under all tasks and settings, the improvement was limited under the OOD setting for OEQA due to training paradigm.
TriviaQA (ID) NQ (OOD)
[…]
Llama2- 7B- Chat Init-Basic 54 46 0 29.3 70.7 0
Init-Refuse 37 21.7 11.5 20.8 38.6 4.8
Van-Tuning 48.6 44.5 -3.7 18.3 50.2 -2.5
Cor-RAIT 41.3 18.3 19.7 16.2 27.6 4.7
CRaFT 38.5 12.9 23.3 15.8 22.4 6.5
Llama3- 8B- Instruct Init-Basic 66.8 33.2 0 40.3 59.7 0
Init-Refuse 53.9 20.8 12 31.1 38.6 5
Van-Tuning 55 38.1 -21.9 21 48.5 -11.7
Cor-RAIT 45.4 13.2 18.8 17.2 25.6 -0.1
CRaFT 43.5 10.9 21.5 19 27.5 0.4
Table 2: Experimental Results in OEQA Task.
MMLU (ID) ARC-c (OOD)
[…]
Llama2- 7B- Chat Init-Basic 45.6 52.8 0 53.9 46 0
Init-Refuse 36.4 38.9 3.9 44.4 35.7 2.6
Van-Tuning 47 53.1 1.2 54.5 45.5 1.2
Cor-RAIT 44.5 39.6 11.3 55.8 38.1 11.1
CRaFT 43.9 36.4 12.5 54.7 35.9 12.6
Llama3- 8B- Instruct Init-Basic 66.8 33.1 0 80.6 19.5 0
Init-Refuse 50 17 15.7 65.3 14.4 5.6
Van-Tuning 69.5 30.5 8 80.3 19.7 -1.3
Cor-RAIT 63.9 21.6 20.4 79.4 16.2 12.2
CRaFT 53.3 9.6 34 74.1 12.7 21.4
Table 3: Experimental Results in MCQA Task.
@@ 6.2 Ablation Experiments
In order to resolve the static and dynamic conflicts that lead to over-refusal, we extend Cor-RAIT to construct RAIT data using the information of correctness, certainty, and knowledge flow. We conduct sufficient ablation experiments to deeply investigate the impact of the above three factors on RAIT data selection. Due to space limitations, only the results of ablation experiments with the LLaMA2-7B-Chat model are reported, as shown in Table 4 and Table 5. Compared to Cor-RAIT, the method only introducing response certainty which named as Cor-RAIT w/o Flow achieved significant gains on the THS in the MCQA and OEQA tasks. This indicates that eliminating static conflicts can effectively mitigate the over-refusal of LLMs and this improvement is generalizable. Cor-RAIT w/o Cer only uses response correctness and knowledge flow. Similarly, experimental results show that introducing knowledge flow to filter dynamic conflicts can also maintain the factuality of the model while improving its rejection capability. Finally, CRaFT considers both static and dynamic conflicts, further enhancing performance improvement.
TriviaQA (ID) NQ (OOD)
[…]
Cor-RAIT 18.32 19.73 27.59 4.71
CRaFT w/o Flow 20.45 21.12 38.75 5.15
CRaFT w/o Cer 15.65 20.11 22.13 5.42
CRaFT 12.93 23.29 22.38 6.54
Table 4: Ablation Experiment Results in OEQA Task.
MMLU (ID) ARC-c (OOD)
[…]
Cor-RAIT 39.57 11.32 38.14 11.08
CRaFT w/o Flow 31.01 12.97 32.34 13.45
CRaFT w/o Cer 32.05 11.51 32.94 13.85
CRaFT 36.39 12.46 35.92 12.58
Table 5: Ablation Experiment Results in MCQA Task
@@ 7 Conclusion
In this paper, we reveal that the over-refusal issue in existing correctness-based RAIT methods stems from the static and dynamic conflicts within the RAIT data. To address these, we introduced CRaFT: Firstly, by incorporating response certainty, we mitigate static conflicts during the construction of RAIT data. Secondly, through rehearsal training, we capture the trends of knowledge flow within the LLM, thereby overcoming dynamic conflicts. We conducted extensive experiments on the MCQA and OEQA tasks. The results show that CRaFT significantly outperforms the existing correctness-based baselines, thereby validating the effectiveness of our method. In the future, we plan to enhance CraFT by integrating RL-based training strategies. We also plan to explore its adaptation for more complex tasks, including reasoning and multi-turn dialogue.
8 Acknowledgments
@@ C.1 Details and Shortcomings of existing refusal-aware metrics
(Xu et al. 2024) defines precision (Prec) as $\frac{N_{c}}{N_{c}+N_{w}}$, where $N_{c}$, $N_{r}$, and $N_{w}$ represent the counts of correct, rejected, and wrong responses, respectively. Accuracy (Acc) is defined as $\frac{N_{c}}{N}$, and truthfulness (Truth) is given by $1-\frac{N_{w}}{N}$. The answer rate (Ans) is $1-\frac{N_{r}}{N}$, representing the proportion of non-rejected responses. The overall reliability (Rely) metric is then:

### v3

@@ 4.1 Overview
Figure 6: The Framework of CRaFT: Stage 1 queries knowledge state and flow, while Stage 2 constructs RAI data and tunes.
@@ 4.2 Query the Knowledge State and Flow of LLM
Correctness is obtained through exact match across the $N$ responses, calculating the proportion of accurate answers. Certainty is evaluated using a pretrained SentenceTransformer model  [^6] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to encode each response $\hat{a}_{j}$ into embedding $E(\hat{a}_{j})$, and the average similarity is computed between these embeddings (excluding diagonal elements). The correctness values range from $[0,1]$. In MCQA task, certainty ranges from [$-log|O|$, 0], and for OEQA, from $[0,1]$. More details about knowledge state query are in Appendix B.1.
@@
4.3 Refusal-Aware Instructions Constuction and Tuning
Unlike Cor-RAIT, which selects RAIT samples solely based on correctness, our approach leverages four parameters $\mu$, $\sigma$, $\Delta\mu$, and $\Delta\sigma$ to characterize both the knowledge state and flow of $\mathcal{M}_{0}$. The challenge lies in making informed sample selections across these four dimensions. We propose a two-step heuristic method outlined in Algorithm 1.
@@ 5.1 Dataset
We evaluate two tasks: knowledge-oriented Multiple Choice Questions Answering (MCQA) and Open-ended Questions Answering (OEQA). For MCQA, the MMLU (Hendrycks et al. 2020) test split serves as the training set, MMLU val as the In-Domain (ID) test set, and ARC-c (Clark et al. 2018) test split as the Out-Of-Domain (OOD) test set. For OEQA, the TriviaQA (Joshi et al. 2017) train split is used for training, TriviaQA dev for the ID test set, and NQ (Kwiatkowski et al. 2019) dev for the OOD test set. More details are in Appendix D.1.
@@ 5.2 Metric
In post RAIT evaluation of LLMs, each test sample is classified as correct, incorrect, or refused. We calculate accuracy ($P_{c}$), error ($P_{w}$), and refusal rates ($P_{r}$) to assess performance, highlighting the key question: How to identify the better-performing model?
@@
We constructed an initial model $\mathcal{M}_{0}$ and four refined models $\mathcal{M}_{1}$ to $\mathcal{M}_{4}$, showing progressive improvement: $\mathcal{M}_{1}<\mathcal{M}_{2}<\mathcal{M}_{3}<\mathcal{M}_{4}$. Details on these models are in Appendix C.2. However, existing metrics have notable flaws: $S_{\text{honesty}}$ (Yang et al. 2023) ranks $\mathcal{M}_{1}$ higher than $\mathcal{M}_{2}$ and treats $\mathcal{M}_{3}$ the same as $\mathcal{M}_{4}$; TRUTHFUL (Cheng et al. 2024) favors $\mathcal{M}_{1}$ over $\mathcal{M}_{2}$; and R-Acc, ER, A-Acc, and A-F1 (Feng et al. 2024) also rank $\mathcal{M}_{1}$ higher than $\mathcal{M}_{2}$. Additionally, AP (Zhang et al. 2024) fails to distinguish between $\mathcal{M}_{3}$ and $\mathcal{M}_{4}$.
@@ 5.3 Baselines
To verify CRaFT’s effectiveness, we compared it with mainstream methods: Init-Basic: Uses the initial LLM with common question-answering prompts. Init-Refuse: Adds instructions like “If you don’t know, respond with ‘I don’t know.’”. Van-Tuning: Randomly selects $N_{\text{van}}+N_{\text{idk}}$ samples from $D_{\text{src}}$ for instruct-tuning without modification. Cor-RAIT: Implements the method from (Zhang et al. 2024), filtering and modifying RAIT data based on response correctness. Detailed prompts for each baseline are in Appendix D.2.
@@ 5.4 Implementation Details
In the experiments, we used LLaMA2-7B-Chat (Touvron et al. 2023) and LLaMA3-8B-Instruct (Dubey et al. 2024) as the initial LLM $\mathcal{M}_{0}$. For the MCQA task, we selected 5000 samples from MMLU, and for the OEQA task, we used 10,000 samples from TriviaQA as training data. In all RAIT settings, except Van-Tuning, the ratio of vanilla to IdK samples was 1:4. We applied 5-shot and 3-shot knowledge state queries for the MCQA and OEQA tasks, respectively. Details on knowledge state and flow queries are in Appendix B. For Instruct Tuning, we used XTuner  [^7] https://github.com/InternLM/xtuner with 3 epochs and a maximum context length of 2048. In MCQA, we applied LoRA (Hu et al. 2021) with settings $r=64$, $\alpha=16$, dropout=0.1, and a learning rate of 2e-4; for OEQA, full parameter training was used. More details on training are in Appendix D.3. We used 0-shot and greedy decoding for evaluation, with further details in Appendix D.4. OpenCompass  [^8] https://github.com/open-compass/opencompass was employed for knowledge state queries and evaluations. All experiments were run on NVIDIA A100-80GB GPUs.
@@ 6 Experimental Results and Analyses
LLMs QA Type MCQA OEQA
Dataset MMLU (ID) ARC-c (OOD) TriviaQA (ID) NQ (OOD)
[…]
LLaMA2-7B Chat Baselines Init-Basic 45.6 52.8 00.0 53.9 46.0 00.0 54.0 46.0 00.0 29.3 70.7 00.0
Init-Refuse 36.4 38.9 03.9 44.4 35.7 02.6 37.0 21.7 11.5 20.8 38.6 04.8
Van-Tuning 46.9 53.1 01.2 54.5 45.5 01.2 48.6 44.5 -03.7 18.3 50.2 -02.5
Cor-RAIT 44.5 39.6 11.3 55.8 38.1 11.1 41.3 18.3 19.7 16.2 27.6 04.7
[…]
w/o Cer 38.4 32.1 11.5 52.5 32.9 13.9 38.5 15.7 20.1 14.6 22.1 05.4
LLaMA3-8B Instruct Baselines Init-Basic 66.8 33.1 00.0 80.6 19.5 00.0 66.8 33.2 00.0 40.3 59.7 00.0
Init-Refuse 50.0 17.0 15.6 65.3 14.4 05.6 53.9 20.8 12.0 31.1 38.6 05.0
Van-Tuning 69.5 30.5 08.0 80.3 19.7 -01.3 55.0 38.1 -21.8 21.0 48.5 -11.7
Cor-RAIT 63.9 21.6 20.4 79.4 16.2 12.2 45.4 13.2 18.8 17.2 25.6 -00.1
[…]
w/o Cer 62.1 18.4 25.0 78.2 17.3 06.5 43.0 11.2 20.5 15.8 23.5 -00.1
Table 2: Performance comparisons on MMLU, ARC-c, TriviaQA and NQ. The best performance is highlighted in boldface.
@@ 6.1 Overall Performance
The experimental results on the OEQA and MCQA tasks are presented in Table 2. Under the ID setting for both types of tasks, our method outperformed four baseline models on THS, achieving the best results. Specifically, under the ID setting for OEQA, compared to the current best RAIT baseline, CRaFT improved the THS on LLaMA2-7B-Chat and LLaMA3-8B-Instruct by 3.56 and 2.72, respectively. Similarly, under the ID setting for MCQA, CRaFT improved the THS by 1.14 and 13.57, respectively. This indicates that CRaFT can significantly improve the model’s rejection capability under the ID setting. Under the OOD setting, CRaFT improved the THS on the MCQA task by 1.5 and 9.2, respectively, compared to Cor-RAIT. On the OEQA’s LLaMA2-7B-Chat, it improved by 1.76 compared to the most competitive method, Init-Refuse. Overall, CRaFT demonstrated excellent competitiveness in model generalization. Furthermore, we found that on the MCQA task, compared to other baselines, Cor-RAIT showed significant improvements under both ID and OOD settings. However, on the OEQA task, Cor-RAIT performed worse than Init-Refuse under the OOD setting. This reveals the limitations of the instruction fine-tuning method. It’s worth mentioning that Van-Tuning generally had a negative impact on the improvement of overall capability, implying that the instruction fine-tuning approach of forcing the model to answer can undermine the model’s inherent rejection capability. Therefore, although CRaFT surpassed Cor-RAIT under all tasks and settings, the improvement was limited under the OOD setting for OEQA due to training paradigm.
@@ 6.2 Ablation Experiments
In order to resolve the static and dynamic conflicts that lead to over-refusal, we extend Cor-RAIT to construct RAIT data using the information of correctness, certainty, and knowledge flow. We conduct sufficient ablation experiments to deeply investigate the impact of the above three factors on RAIT data selection. Compared to Cor-RAIT, the method only introducing response certainty which named as “w/o Flow” achieved significant gains on the THS in the MCQA and OEQA tasks. This indicates that eliminating static conflicts can effectively mitigate the over-refusal of LLMs and this improvement is generalizable. “w/o Cer” only uses response correctness and knowledge flow. Similarly, experimental results show that introducing knowledge flow to filter dynamic conflicts can also maintain the factuality of the model while improving its rejection capability. Finally, CRaFT considers both static and dynamic conflicts, further enhancing performance improvement.
@@ 7 Conclusion
In this paper, we identify over-refusal in correctness-based RAIT methods, caused by static and dynamic conflicts in RAIT data. To address this, we propose CRaFT: it mitigates static conflicts by incorporating response certainty during data construction and overcomes dynamic conflicts through rehearsal training to capture knowledge flow trends in LLMs. Extensive experiments on MCQA and OEQA tasks show CRaFT outperforms existing baselines, validating its effectiveness. Future work includes enhancing CRaFT with RL-based strategies and adapting it for more complex tasks, such as reasoning and multi-turn dialogue.
Acknowledgments
@@ C.1 Details and Shortcomings of existing refusal-aware metrics
(Xu et al. 2024a) defines precision (Prec) as $\frac{N_{c}}{N_{c}+N_{w}}$, where $N_{c}$, $N_{r}$, and $N_{w}$ represent the counts of correct, rejected, and wrong responses, respectively. Accuracy (Acc) is defined as $\frac{N_{c}}{N}$, and truthfulness (Truth) is given by $1-\frac{N_{w}}{N}$. The answer rate (Ans) is $1-\frac{N_{r}}{N}$, representing the proportion of non-rejected responses. The overall reliability (Rely) metric is then: